# DEFAME: Dynamic Evidence-based FAct-checking with Multimodal Experts

### v2

@@ Abstract
The proliferation of disinformation demands reliable and scalable fact-checking solutions. We present Dynamic Evidence-based FAct-checking with Multimodal Experts (DEFAME), a modular, zero-shot MLLM pipeline for open-domain, text-image claim verification. DEFAME operates in a six-stage process, dynamically selecting the tools and search depth to extract and evaluate textual and visual evidence. Unlike prior approaches that are text-only, lack explainability, or rely solely on parametric knowledge, DEFAME performs end-to-end verification, accounting for images in claims and evidence while generating structured, multimodal reports. Evaluation on the popular benchmarks VERITE, AVeriTeC, and MOCHEG shows that DEFAME surpasses all previous methods, establishing itself as the new state-of-the-art fact-checking system for uni- and multimodal fact-checking. Moreover, we introduce a new benchmark, ClaimReview2024+, featuring claims after the knowledge cutoff of GPT-4o, avoiding data leakage. Here, DEFAME drastically outperforms the GPT-4o Chain-of-Thought baseline, demonstrating temporal generalizability and the potential for real-time fact-checking [^2] We released the code and benchmark dataset publicly at:
https://github.com/multimodal-ai-lab/DEFAME
Machine Learning, ICML
@@ 1 Introduction
In recent years, misinformation has been growing in scale and quality (Chen & Shu, 2024) beyond human capacity to fact-check. “Fake news” has evolved from a lighthearted term into a serious global threat (World Economic Forum, 2024). Driven by higher engagement rates on social media and the increase in AI use, misinformation spreads faster, reaches a broader audience, and causes greater harm  (Li & Xie, 2020; Zannettou et al., ; Wang et al., ; Chen & Shu, 2024). Humans perceive multimodal information as more credible (Newman et al., 2012; Hameleers et al., 2020), often interpreting visuals as “evidence” (Greifeneder et al., 2021), making it particularly persuasive. Approximately $80\%$ of the claims checked by professional fact-checkers are multimodal (Dufour et al., 2024)—signifying a strong priority for checking multimodal content.
@@
Unfortunately, Automated Fact-Checking (AFC) systems are mostly text-only (Dmonte et al., 2024; Vykopal et al., 2024). Only a few works venture into visual claim verification (Khaliq et al., 2024; Shao et al., 2023), but none can handle both multimodal claims and multimodal evidence at once. Most multimodal claim verification systems cannot even retrieve the evidence needed to verify a claim (Fu et al., 2024; Vo-Hoang et al., 2024; Tang et al., 2024),. The majority of AFC works focuses on only a specific aspect of AFC such as evidence retrieval (Cheung & Lam, 2023), evidence summarization (Chen et al., 2024), or evidence ranking (Tahmasebi et al., 2024). This specialization has created a scattered landscape where individual approaches address only isolated aspects of a complex problem. AFC systems that target the overall task of fact-checking typically are text-only (Zhao et al., 2024; Li et al., 2024c), lack performance (Chen et al., 2024; Cao et al., 2023; Papadopoulos et al., 2024a), do not involve a flexible planning module (Khaliq et al., 2024; Tahmasebi et al., 2024), or are not explainable (Shao et al., 2023; Xu et al., 2024).
Therefore, we introduce DEFAME, a straightforward end-to-end AFC framework, unifying the advancements in the field of AFC into one single system. As the first of its kind, it is able to natively process multimodal claims and evidence, the latter of which it retrieves dynamically as needed. DEFAME is designed with transparency in mind, imitating a human fact-checking process and returning a detailed fact-check report to the user (cf. Figure 1).
Our contributions can be summarized as follows: First, we present DEFAME, the first multimodal AFC system that can handle multimodal claims as well as retrieve and process multimodal evidence. DEFAME brings together the advancements of previous work by unifying their strengths into one single, yet straightforward framework.
Second, we empirically demonstrate DEFAME’s effectiveness by establishing new state-of-the-art results on three diverse and widely used benchmarks. On AVeriTeC (Schlichtkrull et al., 2024b), we improve accuracy from $65.6\%$ to $74.0\%$; on MOCHEG (Yao et al., 2023), we achieve a $+10.9\%$ improvement in micro F1 score, and on VERITE (Papadopoulos et al., 2024b), we enhance True/False accuracy by $+26.5\%$.
Third, we contribute a new benchmark, ClaimReview2024+, with claims that occurred after the knowledge cutoff of GPT-4o to mitigate the effects of data leakage and show that DEFAME is strongly superior to GPT-4o on these “unseen” statements.
Fourth, we demonstrate that the fact-checking reports generated by DEFAME are well-received by human evaluators, preferring the reports over GPT-4o’s outputs.
Finally, we made our code and benchmark publicly available [^3] https://github.com/multimodal-ai-lab/DEFAME.
@@ 2 Related Work
Method Multimodal claims Multimodal evidence Evidence retrieval Multi-hop Planning Tools Reasoning Explainable Open- domain Backbone
RAGAR (Khaliq et al., 2024) - - - - MLLM
MMOOC-Checker (Xu et al., 2024) - - - - - - - VLT
Vo-Hoang et al. (Vo-Hoang et al., 2024) - - - - - - - LLM & VLT
Zeng et al. (Zeng et al., 2024a) - - - - - - - MLLM
CHASMA (Papadopoulos et al., 2024b) - - - - - - - CLIP & VLT
AITR (Papadopoulos et al., 2024a) - - - - - - - VLT
HAMMER (Shao et al., 2023) - - - - - - - - BERT & ViT
MultiMD (Fu et al., 2024) - - - - - - - - VLT
@@
MOCHEG (Yao et al., 2023) - - - - - VLT
MetaSum (Chen et al., 2024) - - - - - VLT & LLM
M³D (Tang et al., 2024) - - - - - - - - VLT & GCN
ChartBERT (Akhtar et al., 2023a) - - - - - - - - BERT
PACAR (Zhao et al., 2024) - - LLM
HiSS (Zhang & Gao, 2023) - - - - LLM
@@
Self-Checker (Li et al., 2024c) - - - - LLM
FactLLaMA (Cheung & Lam, 2023) - - - - - - - LLM
CFR (Sriram et al., 2024) - - - - - - - BERT & LLM
GPT-CoT (Cao et al., 2023) - - - - - - LLM
@@
Automating the task of fact-checking is a difficult problem, far from being solved (Akhtar et al., 2023b; Dmonte et al., 2024; Yao et al., 2023; Schlichtkrull et al., 2024a). Due to its complexity, Akhtar et al. (2023b) subdivide AFC into three stages: (1) claim detection & extraction, (2) evidence retrieval, and (3) verdict prediction. Most approaches narrow down their scope, either by focusing on a sub-task like summarization (Chen et al., 2024), justification generation  (Atanasova et al., 2020), evidence retrieval  (Samarinas et al., 2021), deepfake detection (Jia et al., 2024), out-of-context detection (Xu et al., 2024; Vo-Hoang et al., 2024), or by addressing only a specific domain, e.g., charts (Akhtar et al., 2023a, 2024), social media (Wang et al., 2018), politics (Khaliq et al., 2024), or news (Xu et al., 2024; Shao et al., 2023). Others investigate the incorporation of new architectures like a knowledge graph (Cao et al., 2024) or address the problem of evidence ambiguity (Glockner et al., 2024). In contrast to all these methods, DEFAME combines the fragmented landscape of AFC work into one end-to-end solution—not restricted to only one modality, domain, or sub-task. Table 1 compares DEFAME to prior work.
Text-only fact-checking
@@
The vast majority of proposed AFC systems is purely text-based (Hassan et al., 2017; Thorne et al., 2018; Schlichtkrull et al., 2024b; Dmonte et al., 2024; Vykopal et al., 2024; Yang & Rocha, 2024; Zhao et al., 2024; Wang et al., 2024b; Li et al., 2024b; Pan et al., 2023; Schlichtkrull et al., 2024a; Cheung & Lam, 2023). The most popular benchmarks to evaluate these methods are LIAR (Wang, 2017), FEVER (Thorne et al., 2018), and AVeriTeC (Schlichtkrull et al., 2024b), the latter of which mitigates important weaknesses of the previous ones, including reliance on artificial claims and evidence insufficiency. For this reason, we use AVeriTeC in our evaluation. Most recently, Factcheck-Bench (Wang et al., 2024a) was introduced to evaluate the factuality of entire LLM responses.
Multimodal fact-checking
@@
Popular multimodal AFC benchmarks include works from Xu et al. (2024); Nielsen & McConville (2022); Zlatkova et al. (2019); Nakamura et al. (2020); Aneja et al. (2021); Yao et al. (2023); Jaiswal et al. (2017); Sabir et al. (2018); Müller-Budack et al. (2020); Luo et al. (2021a). Notably, MOCHEG (Yao et al., 2023) builds on real-world claims, requiring multimodal evidence retrieval, additionally incorporating the task of justification generation. Given these features, we use MOCHEG in our evaluation. A strong emphasis of multimodal AFC has been on Out-Of-Context (OOC) image detection, sometimes referred to as “cheapfake” detection. Popular evaluation benchmarks include NewsCLIPpings (Luo et al., 2021b) and, more recently, VERITE (Papadopoulos et al., 2024b). In our experiments, we use VERITE because it improves over previous benchmarks by reducing unimodal bias and incorporating real-world samples.
@@
(M)LLM-based fact-checking
With the rise of (Multimodal) Large Language Models or (M)LLMs, the AFC community increasingly explored prompt-driven solutions (Yang & Rocha, 2024; Khaliq et al., 2024; Tahmasebi et al., 2024; Chen et al., 2024; Pan et al., 2023; Cheung & Lam, 2023). One of DEFAME’s closest relatives, RAGAR (Khaliq et al., 2024), processes textual and visual claims but retrieves only textual evidence. Furthermore, DEFAME directly incorporates the claim image in its context while RAGAR converts it into a textual description, discarding critical visual information.
While there have been efforts to improve the performance of MLLM-based approaches through training on synthetic data  (Zeng et al., 2024a), most still rely on the parametric knowledge of the MLLMs (Beigi et al., 2024; Shao et al., 2023; Li et al., 2024e; Geng et al., 2024), foregoing external evidence retrieval. This approach has three major drawbacks: (1) MLLM knowledge is static and fails on recent claims, as shown in this work; (2) predictions lack links to verifiable sources, reducing transparency; and (3) reliance on parametric knowledge increases hallucination risks, making such methods less reliable than Retrieval-Augmented Generation (RAG)-based approaches. In contrast, DEFAME uses internal knowledge only for commonsense reasoning, retrieving all evidence dynamically through external tools. Unlike some prior work (Yang & Rocha, 2024; Tang et al., 2024), it does not rely on benchmark-provided gold evidence, reinforcing its adaptability to real-world misinformation.
3 DEFAME Approach
@@
Stage 1: Plan Actions
@@
Stage 2: Execute Actions
@@
To prevent temporal leakage, all web-based tools restrict search results to sources published before the claim’s release date (if known). Additionally, we exclude major fact-checking websites and any sites that disallow automated bot access. A list of excluded domains can be found in Appendix A. For each retrieved URL, we scrape the corresponding page using Firecrawl [^6] https://github.com/mendableai/firecrawl. Unlike previous work, we extend the scraper to identify and download any referenced image, ensuring a complete context for the fact-check.
Stage 3: Summarize Results
@@
Stage 4: Develop the Fact-Check
@@
Corresponding to stage 4 in Moreno Gil et al. (2021), DEFAME brings claim and summarized evidence together. It directs the MLLM to discuss the claim’s veracity step-by-step based on the evidence, flagging any gaps as “incomplete” if the information is missing. This stage offers room for intricate reasoning, deducing new insights through natural language inference.
Stage 5: Predict a Verdict
At this stage, DEFAME assesses the claim’s veracity, classifying it into one of the benchmark-specific categories. The MLLM is prompted to summarize key findings from the report and select a verdict. If there is insufficient information, it may return [Uncaptioned image]NEI. When [Uncaptioned image]NEI is returned, the fact-checking process returns to Stage 1 to gather further evidence. This allows DEFAME to dynamically dive deeper into unresolved questions, allowing for a more thorough exploration of relevant sub-branches. DEFAME moves to the final stage if the end of the third iteration or a definitive verdict is reached, mirroring the systematic approach of human fact-checkers.
Stage 6: Justify the Verdict
Human readability and explainability are central in this final stage. Given the entire fact-checking report, DEFAME tasks the MLLM with generating a concise summary, highlighting key findings and critical evidence—including hyperlinks and images. The final output transparently documents the entire process, aiding verification and further human fact-checking.

### v3

@@ Abstract
The proliferation of disinformation demands reliable and scalable fact-checking solutions. We present Dynamic Evidence-based FAct-checking with Multimodal Experts (DEFAME), a modular, zero-shot MLLM pipeline for open-domain, text-image claim verification. DEFAME operates in a six-stage process, dynamically selecting the tools and search depth to extract and evaluate textual and visual evidence. Unlike prior approaches that are text-only, lack explainability, or rely solely on parametric knowledge, DEFAME performs end-to-end verification, accounting for images in claims and evidence while generating structured, multimodal reports. Evaluation on the popular benchmarks VERITE, AVeriTeC, and MOCHEG shows that DEFAME surpasses all previous methods, establishing itself as the new general state-of-the-art fact-checking system for uni- and multimodal fact-checking. Moreover, we introduce a new multimodal benchmark, ClaimReview2024+, featuring claims after the knowledge cutoff of GPT-4o, avoiding data leakage. Here, DEFAME drastically outperforms the GPT-4o baselines, showing temporal generalizability and the potential for real-time fact-checking [^2] We released the code and benchmark dataset publicly at:
https://github.com/multimodal-ai-lab/DEFAME/tree/icml
Machine Learning, ICML, Fact-Checking, Claim Verification, Large Language Models, Multimodal
@@ 1 Introduction
In recent years, misinformation has been growing in scale and quality (Chen & Shu, 2024) beyond human capacity to fact-check. “Fake news” has evolved from a lighthearted term into a serious global threat (World Economic Forum, 2024). Driven by higher engagement rates on social media and the increase in AI use, misinformation spreads faster, reaches a broader audience, and causes greater harm  (Li & Xie, 2020; Zannettou et al., 2018; Wang et al., 2020; Chen & Shu, 2024). Humans perceive multimodal information as more credible (Newman et al., 2012; Hameleers et al., 2020), often interpreting visuals as “evidence” (Greifeneder et al., 2021), making it particularly persuasive. Approximately $80\%$ of the claims checked by professional fact-checkers are multimodal (Dufour et al., 2024)—signifying a strong priority for checking multimodal content.
@@
Unfortunately, Automated Fact-Checking (AFC) systems are mostly text-only (Dmonte et al., 2024; Vykopal et al., 2024). Only a few works venture into visual claim verification (Khaliq et al., 2024; Shao et al., 2023), but none can handle both multimodal claims and multimodal evidence at once. Most multimodal claim verification systems cannot retrieve the evidence needed to verify a claim (Fu et al., 2024; Vo-Hoang et al., 2024; Tang et al., 2024). The majority of AFC works focus on a specific aspect of AFC such as evidence retrieval (Cheung & Lam, 2023), evidence summarization (Chen et al., 2024), or evidence ranking (Tahmasebi et al., 2024). This specialization has created a scattered landscape where individual approaches address isolated aspects of a complex problem. AFC systems that target the overall task of fact-checking typically are text-only (Zhao et al., 2024; Li et al., 2024c), lack performance (Chen et al., 2024; Cao et al., 2023; Papadopoulos et al., 2024a), do not involve a flexible planning module (Khaliq et al., 2024; Tahmasebi et al., 2024), or are not explainable (Shao et al., 2023; Xu et al., 2024).
Therefore, as the main contribution of this paper, we introduce DEFAME, a straightforward end-to-end AFC framework, unifying the advancements in the field of AFC into one single system. As the first of its kind, it is able to natively process multimodal claims and evidence, the latter of which it retrieves dynamically as needed. DEFAME is designed with transparency in mind, imitating a human fact-checking process and returning a detailed fact-check report to the user (cf. Figure 1).
We empirically demonstrate DEFAME’s effectiveness and generality by establishing new state-of-the-art results on three diverse and widely used benchmarks, surpassing even the most specialized methods. On AVeriTeC (Schlichtkrull et al., 2024b), we improve accuracy from $65.6\%$ to $70.5\%$; on MOCHEG (Yao et al., 2023), we achieve a $+10.6\%$ improvement in accuracy, and on VERITE (Papadopoulos et al., 2024b), we enhance True/False accuracy by $+25.9\%$.
Additionally, we contribute a new benchmark, ClaimReview2024+, with claims that occurred after the knowledge cutoff of GPT-4o to mitigate the effects of data leakage. We show that DEFAME is superior to GPT-4o on these “unseen” statements. We show that the fact-checking reports generated by DEFAME are well-received by human evaluators, who prefer them over GPT-4o’s outputs. Finally, we make our code and benchmark publicly available [^3] https://github.com/multimodal-ai-lab/DEFAME/tree/icml.
@@ 2 Related Work
Method     Multimodal claims     Multimodal evid.     Evidence retrieval     Multi-hop     Planning     Tools     Reasoning     Explainable     Training-free     Open-domain     Open source Backbone
RAGAR (Khaliq et al., 2024) - - - - - MLLM
MMD-Agent (Liu et al., 2025) - - - - MLLM
SNIFFER (Qi et al., 2024) - - - - - VLT & LLM
MMOOC-Checker (Xu et al., 2024) - - - - - - - - - VLT
Vo-Hoang et al. (Vo-Hoang et al., 2024) - - - - - - - - - LLM & VLT
Zeng et al. (Zeng et al., 2024a) - - - - - - - - - MLLM
CHASMA (Papadopoulos et al., 2024b) - - - - - - - - CLIP & VLT
AITR (Papadopoulos et al., 2024a) - - - - - - - - VLT
HAMMER (Shao et al., 2023) - - - - - - - - - BERT & ViT
MultiMD (Fu et al., 2024) - - - - - - - - - - VLT
@@
MOCHEG (Yao et al., 2023) - - - - - - VLT
MetaSum (Chen et al., 2024) - - - - - - - VLT & LLM
M³D (Tang et al., 2024) - - - - - - - - - - VLT & GCN
ChartBERT (Akhtar et al., 2023a) - - - - - - - - - - BERT
PACAR (Zhao et al., 2024) - - - LLM
HiSS (Zhang & Gao, 2023) - - - - - LLM
@@
Self-Checker (Li et al., 2024c) - - - - - LLM
FactLLaMA (Cheung & Lam, 2023) - - - - - - - - - LLM
CFR (Sriram et al., 2024) - - - - - - - - - BERT & LLM
DeBERTa (Cao et al., 2023) - - - - - - - LLM
@@
Automating the task of fact-checking is a difficult problem, far from being solved (Akhtar et al., 2023b; Dmonte et al., 2024; Yao et al., 2023; Schlichtkrull et al., 2024a). Due to its complexity, Akhtar et al. (2023b) subdivide AFC into three stages: (1) claim detection & extraction, (2) evidence retrieval, and (3) verdict prediction. Most approaches narrow down their scope, either by focusing on a sub-task like summarization (Chen et al., 2024), justification generation  (Atanasova et al., 2020), evidence retrieval  (Samarinas et al., 2021), deepfake detection (Jia et al., 2024), out-of-context detection (Xu et al., 2024; Vo-Hoang et al., 2024), image contextualization (Tonglet et al., 2024), or by addressing only a specific domain, e.g., charts (Akhtar et al., 2023a, 2024), social media (Wang et al., 2018), politics (Khaliq et al., 2024), or news (Xu et al., 2024; Shao et al., 2023). Others investigate the incorporation of new architectures like a knowledge graph (Cao et al., 2024) or address the problem of evidence ambiguity (Glockner et al., 2024). In contrast to all these methods, DEFAME combines the fragmented landscape of AFC work into one end-to-end solution—not restricted to only one modality, domain, or sub-task. Table 1 compares DEFAME to prior work.
@@
Text-only fact-checking: The vast majority of proposed AFC systems is purely text-based (Hassan et al., 2017; Thorne et al., 2018; Schlichtkrull et al., 2024b; Dmonte et al., 2024; Vykopal et al., 2024; Yang & Rocha, 2024; Zhao et al., 2024; Wang et al., 2024b; Li et al., 2024b; Pan et al., 2023; Schlichtkrull et al., 2024a; Cheung & Lam, 2023). The most popular benchmarks to evaluate these methods are LIAR (Wang, 2017), FEVER (Thorne et al., 2018), and AVeriTeC (Schlichtkrull et al., 2024b), the latter of which mitigates important weaknesses of the previous ones, including reliance on artificial claims and evidence insufficiency. For this reason, we use AVeriTeC in our evaluation. Most recently, Factcheck-Bench (Wang et al., 2024a) was introduced to evaluate the factuality of entire LLM responses.
@@
Multimodal fact-checking: Popular multimodal AFC benchmarks include works from Xu et al. (2024); Nielsen & McConville (2022); Zlatkova et al. (2019); Nakamura et al. (2020); Aneja et al. (2021); Yao et al. (2023); Jaiswal et al. (2017); Sabir et al. (2018); Müller-Budack et al. (2020); Luo et al. (2021a). Notably, MOCHEG (Yao et al., 2023) builds on real-world claims, requiring multimodal evidence retrieval, additionally incorporating the task of justification generation. Given these features, we use MOCHEG in our evaluation. A strong emphasis of multimodal AFC has been on Out-Of-Context (OOC) image detection, sometimes referred to as “cheapfake” detection. Popular evaluation benchmarks include NewsCLIPpings (Luo et al., 2021b) and, more recently, VERITE (Papadopoulos et al., 2024b). In our experiments, we use VERITE because it improves over previous benchmarks by reducing unimodal bias and incorporating real-world samples.
@@
(M)LLM-based fact-checking: With the rise of (Multimodal) Large Language Models or (M)LLMs, the AFC community increasingly explored prompt-driven solutions (Yang & Rocha, 2024; Khaliq et al., 2024; Tahmasebi et al., 2024; Chen et al., 2024; Pan et al., 2023; Cheung & Lam, 2023). One of DEFAME’s closest relatives, RAGAR (Khaliq et al., 2024), processes textual and visual claims but retrieves only textual evidence. Furthermore, DEFAME directly incorporates the claim image in its context while RAGAR converts it into a textual description, discarding critical visual information.
While there have been efforts to improve the performance of MLLM-based approaches through training on synthetic data  (Zeng et al., 2024a), most still rely on the parametric knowledge of the MLLMs (Beigi et al., 2024; Shao et al., 2023; Li et al., 2024e; Geng et al., 2024), foregoing external evidence retrieval. This approach has three major drawbacks: (1) MLLM knowledge is static and fails on recent claims, as shown in this work; (2) predictions lack links to verifiable sources, reducing transparency; and (3) reliance on parametric knowledge increases hallucination risks, making such methods less reliable than Retrieval-Augmented Generation (RAG)-based approaches. In contrast, DEFAME uses internal knowledge mostly for commonsense reasoning, retrieving evidence dynamically through external tools. Unlike some prior work (Yang & Rocha, 2024; Tang et al., 2024), it does not rely on benchmark-provided gold evidence, reinforcing its adaptability to real-world misinformation.
3 DEFAME Approach.
@@
Stage 1: Plan Actions.
@@
Stage 2: Execute Actions.
@@
To prevent temporal leakage, all web-based tools restrict search results to sources published before the claim’s release date (if known). Additionally, we exclude major fact-checking websites and any sites that disallow automated bot access. A list of excluded domains can be found in Appendix D. For each retrieved URL, we scrape the corresponding page using Firecrawl [^6] https://github.com/mendableai/firecrawl. Unlike previous work, we extend the scraper to identify and download any referenced image, ensuring a complete context for the fact-check.
Stage 3: Summarize Results.
@@
Stage 4: Develop the Fact-Check.
@@
Corresponding to Stage 4 in Moreno Gil et al. (2021), DEFAME brings claim and summarized evidence together. It directs the MLLM to discuss the claim’s veracity step-by-step based on the evidence, flagging any gaps as “incomplete” if the information is missing. This stage offers room for intricate reasoning, deducing new insights through natural language inference. It serves as an immediate preparation for the following stage.
Stage 5: Predict a Verdict.
Next, DEFAME classifies the claim into one of the benchmark-specific categories by prompting the MLLM to summarize key findings and select a verdict. If the model returns [Uncaptioned image]NEI (Not Enough Information), the system loops back to Stage 1 to retrieve additional evidence, enabling deeper exploration of unresolved aspects. The process proceeds to the final stage once a definitive verdict is reached or three iterations have been completed, reflecting the iterative nature of human fact-checking.
Stage 6: Justify the Verdict.
Human readability and explainability are essential to effective fact-checking. While the fact-checking report provides a detailed and transparent account of the verification process, its length can become very long—potentially overwhelming users seeking a quick understanding of the outcome. This final, post-prediction stage addresses that need by generating a concise summary that distills the key findings and critical evidence, including hyperlinks. To this end, the MLLM is prompted with the complete report and tasked with producing a focused, readable justification. The resulting summary is appended to the full report, serving both as an accessible explanation for end users and as a support tool for further human verification.
4 ClaimReview2024+
We introduce ClaimReview2024+, a novel benchmark designed to evaluate fact-checking models on claims that fall outside the pretraining scope of current MLLMs, avoiding the data leakage issue (see Appendix A for more details). To this end, we curate a set of $300$ English claims from the Google FactCheck Claim Search API [^7] https://toolbox.google.com/factcheck/apis, which indexes professionally fact-checked claims published by organizations such as Snopes, PolitiFact, and AFP via the ClaimReview markup standard [^8] https://www.claimreviewproject.com/. All claims are sampled from fact-checking articles from the period between November 1, 2023, and January 18, 2025—notably, after the October 2023 knowledge cutoff of GPT-4o—enabling us to simulate a realistic, temporally out-of-distribution evaluation setting. The dataset consists of $160$ unimodal (text-only) and $140$ multimodal (text-image) claims (cf. Fig. 3).
Figure 3: Examples from our ClaimReview2024+ benchmark, containing claimant, claim date, claim, and verdict.
Claims are retrieved via keyword-based queries across ten broad topics (e.g., politics, health, climate) and de-duplicated based on review URLs. For each instance, we extract the date and the claimant for a clearer claim context. If the claim date is not present in the markup, we use the review’s release date instead. Labels are adapted from those provided by the fact-checking organizations mapped to a unified four-class taxonomy inspired by AVeriTeC: [Uncaptioned image]Supported, [Uncaptioned image]Refuted, [Uncaptioned image]Misleading, and [Uncaptioned image]NEI—see Appendix B for the specific class definitions. Label assignment follows a two-stage process: Trivial cases are mapped automatically using an LLM-based script (included in the release), while ambiguous ones are manually labeled by a PhD-level AFC researcher. The dataset is validated by a second annotator, comparing all instances with the original fact-checking articles directly.
To prevent leakage of veracity judgments, we manually edit claims that contain explicit fact-checking language (e.g., rewriting “Fact Check: Image from Pakistan falsely claims…” to “Image from Pakistan shows…”). Since the ClaimReview API often returns teaser images with overlays or composites, all images are manually curated to reflect the original visuals referenced in the claims.
The final label distribution is: $129$ [Uncaptioned image]Refuted, $89$ [Uncaptioned image]Supported, $61$ [Uncaptioned image]Misleading, and $21$ [Uncaptioned image]NEI. While we ensure all claims are posted after the GPT-4o cutoff date, we note that some may represent earlier events re-evaluated by fact-checking organizations. Given its source diversity and topical focus on timely, high-impact misinformation, we believe ClaimReview2024+ offers a challenging, real-world testbed for current fact-checking models.